OpenStudy (anonymous):

~I will fan and medal~ Part B: Solve 9x2 + 3x − 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (5 points) Please answer ASAP!

OpenStudy (danjs):

Cant go wrong always using the quadratic formula

OpenStudy (danjs):

i would try to factor it first before that though...

OpenStudy (anonymous):

True.. I'll try that. @DanJS

OpenStudy (danjs):

9 1 2 1 3 3 i think you can make that +3 in the middle from those

OpenStudy (danjs):

3*2 - 3*1

OpenStudy (danjs):

Or just use the quadratic formula,, reason: always works.. :)

OpenStudy (anonymous):

When I used the quadratic formula, I got 0.3333333~ and -0.66666666666667..

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

hmm... real fast by factoring first \[9x^2+3x-2 = (3x - 1)*(3x+2) = 0\]

OpenStudy (danjs):

so either of those quantities in ( ) can be zero to be true 3x - 1 = 0 or 3x + 2 = 0

OpenStudy (danjs):

so you have it right, just decimal form

OpenStudy (danjs):

x = + 1/3 or x = - 2/3 i would leave it like that

OpenStudy (anonymous):

Alright, I got it now. Thanks for the help!

OpenStudy (danjs):

welcome, anytime